DIFFERENTIAL GEOMETRY/PDE SEMINAR

Wednesday, May 2 Communications 232 3:45 PM

Higher dimensional minimal hypersurfaces

Frank Pacard (Université Paris XII and MSRI)

Many procedures are now available to produce interesting minimal surfaces in R^{3}. In higher dimensions not so many examples of minimal hypersurfaces are known.

In all dimensions, we will describe a reasonnable generalization of three families of minimal surfaces: the k-noids surfaces (which are complete minimal surfaces with catenoidal ends), the Riemann’s surface family (which is a one parameter family of singly periodic minimal surfaces folliated horizontally by circles) and finaly the Scherk’s second surface family (which is a one parameter family of singly periodic minimal surfaces with 4 planar ends).